OpenStudy (anonymous):

Please show me how to solve this? F(x) = 1/x-5 and g(x) = x^2+2 Find: A... (f o g)(x) B... (g o f)(6)

hartnn (hartnn):

to get (fog)(x) just replace x in f(x) by g(x)

OpenStudy (anonymous):

You're going to replace the x in f(x) with g(x), so it'll be f(g(x)) = 1/(x^2+2)-5.

OpenStudy (anonymous):

So the answer would be 1/-3... ?

hartnn (hartnn):

where did u get 1/-3 from ?

OpenStudy (anonymous):

1 all over (x^2+2)-5 x^2+2 = 2 - 5 = -3

hartnn (hartnn):

why x^2+2 = 2 - 5 = -3 ?? (fog) (x) is just 1 all over( (x^2+2)-5) ok ?

OpenStudy (anonymous):

I don't solve it any further?

hartnn (hartnn):

then u get 1 all over( (x^2+2)-5) = 1 all over(x^2-3)

OpenStudy (anonymous):

Ohhhh. Ok, that makes sense.

hartnn (hartnn):

now find (gof)(x)

OpenStudy (anonymous):

Do I take 1 all over x^2-3 and replace the x with 6? Or do I start with the original equations?

hartnn (hartnn):

to find (gof)(x) replace all 'x' in g(x) by f(x) = 1/(x-5)

hartnn (hartnn):

so start with original equations

OpenStudy (anonymous):

Sooo. g(6) = 1 all over 6-5 = 1 then g(1) = 1^2+2 = 3

hartnn (hartnn):

that is correct, u just made one typo: f(6) = 1 all over 6-5 = 1 then g(1) = 1^2+2 = 3

hartnn (hartnn):

so (gof)(6)=3

OpenStudy (anonymous):

Thaaaaank you so much!

hartnn (hartnn):

welcome :)